OpenStudy (anonymous):

Soybean meal is 16% protein. How many pounds of each should be mixed together to get 320 lb mixture that is 15% protein? How many lbs of cornmeal? How many lbs of soybean?

OpenStudy (dumbcow):

What percent of cornmeal is protein?

OpenStudy (anonymous):

Cornmeal is 8 %

OpenStudy (dumbcow):

Try setting up a table: Lbs Percent Protein Soybean: S 16 16*S Corn: C 8 8*C -------------------------- Total: 32 15 32*15 where 1st and 3rd columns should add up to total S+C = 32 16S +8C = 32*15

OpenStudy (dumbcow):

sorry, 320 not 32

OpenStudy (dumbcow):

Then use elimination or substitution to solve for S and C

OpenStudy (dumbcow):

S+C = 320 16S +8C = 4800 Multiply top equation by -8, to eliminate variable C -8S -8C = -2560 16S+8C = 4800 ---------------- 8S = 2240 S = 280 Substitute it back in to solve for C